OpenStudy (anonymous):

3r = (dr/dtheta) - theta^3. Determine if it's linear, seperable, neither, or both

OpenStudy (zzr0ck3r):

\[3r=\frac{dr}{d\theta}-\theta^3\]

OpenStudy (zzr0ck3r):

can we get all the r,dr's on one side and theta,dthetas on the other?

OpenStudy (anonymous):

So this is \[3r=\frac{dr}{d\theta}-\theta^3\] I will replace the substitution: \(r=y\) \(\theta=x\) Now we have: \[3y=\frac{dy}{dx}-x^3\] We can re-arrange this to: \[\frac{dy}{dx}-3y=x^3\] This is in the form of: \[\frac{dy}{dx}+p(x)y=q(x)\] So therefore, it is linear and we can separate it

OpenStudy (anonymous):

We isolate differential equations of the form: \[\frac{dy}{dx}+p(x)y=q(x)\] By using this formula: \[y=\frac{\int{e^{p(x)dx}q(x)}+C}{e^{p(x) dx}}\]

OpenStudy (anonymous):

This makes so much more since.

OpenStudy (anonymous):

Or simplified: \[y=\frac{\int{\mu(x)q(x)}+C}{\mu(x)}\] Where \(\mu(x)=e^{\int{p(x)dx}}\) And I made a mistake on the above one

OpenStudy (anonymous):

So we can simplify this by stating: \[\eqalign{ &p(x)=-3 \\ &q(x)=x^3 \\ }\] So: \[y=\frac{\int{\mu(x)x^3}+c_1}{\mu(x)}\] Where: \[\mu(x)=e^{\int{-3 dx}}=e^{-3x}\] So therefore: \[y=\frac{\int{e^{-3x}x^3}+c_1}{e^{-3x}}\]

OpenStudy (anonymous):

Very thorough explanation. Thank you

OpenStudy (anonymous):

Your welcome :) I could go a step further and simplify evaluating the integrals if you want. Or I can stop there if you're happy enough.

OpenStudy (anonymous):

This is fine although I do need help with another question if you're interested.